### v1

# Standard modules of affine Hecke algebras

Abstract
Let $G$ be a connected reductive group defined and split over a non-archimedean local field $F$. We give a new geometric proof of a special case of a recent theorem of Solleveld. Namely, we show that the class of standard Iwahori-spherical $G(F)$-representations, a notion a priori dependent on the coefficient field being the complex numbers, is actually defined over $\bar{\mathbb{Q}}_{\ell}$.
An unpublished theorem of Clozel, proven with global techniques, says that the class of essentially square-integrable representations is also defined over $\bar{\mathbb{Q}}_{\ell}$. As an application of our main result, we give a local proof of this theorem for $G=\mathrm{GL}_{n}$.
Keywords— Standard representation, essentially square-integrable representation, Iwahori-Hecke algebra, Chow group
1 Introduction
1.1 Representations over $\bar{\mathbb{Q}}_{\ell}$
Let $G$ be a connected reductive group defined and split over a non-archimedean local field $F$. Let $\pi$ be a complex admissible representation of $G(F)$, let $\gamma$ be a field automorphism of $\mathbb{C}$, and define a representation $\gamma\cdot\pi$ of $G(F)$ by declaring that its matrix coefficients are of the form $\gamma(f(g))$, for matrix coefficients $f$ of $\pi$ (see Section 2.2 for a precise definition). It is clear that $\gamma\cdot\pi$ is irreducible when $\pi$ is, and likewise for supercuspidality. In [KSV] and [SolleveldStandard], less obvious heritability from $\pi$ to $\gamma\cdot\pi$ is studied. Recall that the standard representations of $G(F)$ are by definition parabolic inductions $i_{P}^{G}(\sigma\otimes\nu)$ for irreducible tempered representations $\sigma$ and positive unramified characters $\nu$ of of $M_{P}(F)$, respectively, where $M_{P}$ is the Levi subgroup of $P$ (see [SolleveldStandard, Def. 4.1, Thm. 4.3 (a)]). Thus the standard modules occur in algebraic families.
As pointed out in [SolleveldStandard], this occurrence in families is one reason to think that the standard representations may be better behaved in the context of the categorical Local Langlands correspondence of [FS] than simple representations. For this to make sense, though, the notion of “standard representation” should be defined over $\bar{\mathbb{Q}}_{\ell}$, i.e., be stable under the above action of $\mathrm{Aut}(\mathbb{C})$. To this end, Solleveld has recently shown
Theorem 1 ([SolleveldStandard], Prop. 5.5).
Let $G$ be a connected reductive group over $F$ and $\pi$ a standard representation of $G(F)$. Then $\gamma\cdot\pi$ is also a standard representation of $G(F)$.
The first result of this note is a new geometric proof, for representations in the principal block of split groups, of Theorem 1. Precisely, we give a geometric proof, as our Corollary 1, that if $K(s,N,\rho)$ is a standard module for the Iwahori-Hecke algebra $H$ of $G$, as constructed in terms of $G^{\vee}$ and the enhanced $L$-parameter $(s,N,\rho)$ in [KLDeligneLanglands], [CG], then $\gamma\cdot K(s,N,\rho)\simeq K(\gamma(s),N,\rho_{\gamma})$, as one would expect upon considering central characters. Here, $\rho_{\gamma}$ is the pullback of $\rho$ along the isomorphism $\pi_{0}(Z_{G^{\vee}}(s,N))\simeq\pi_{0}(Z_{G^{\vee}}(\gamma(s),N))$ induced by $\gamma$, see (10) below.
Note that if $\pi$ is generated by its Iwahori-fixed vectors for a chosen Iwahori subgroup $I$ of $G(F)$, then so is $\gamma\cdot\pi$. In general, the functor $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}$- does not preserve Bernstein blocks (consider $G=\mathrm{GL}_{1}$), but as it is an auto-equivalence of the category of smooth representations, it still interacts well enough with types and Hecke algebras for our purposes, see Section 3.1.
The main input to our argument is the main result of [DLP], which with [KLDeligneLanglands], [CG] implies that, up the central character, representations $K(s,N,\rho)$ can be constructed on rationalized Chow groups of certain varieties defined over $\mathbb{Q}$. This may be viewed as further suggestion that (at least for the principal block) the standard modules admit a natural interpretation in terms of [FS].
After to appealing to [DLP], the argument is very formal, amounting essentially to various intersection-theoretic constructions commuting with isomorphisms of schemes, i.e., being well-defined. Nevertheless, we have tried to include careful references for these constructions beyond the usual references for singular homology, because it is conceptually important that we may work morphisms of schemes that are not morphisms of $\mathbb{C}$-schemes.
Recall that a Langlands parameter
$W_{F}\times\mathrm{SL}_{2}\to G^{\vee}(\mathbb{C})$
is called essentially discrete if it does not factor through any proper Levi subgroup of $G^{\vee}$. Clearly, this property is stable under post-composition with field automorphisms of $\mathbb{C}$. The corresponding property for representations is of course not obvious. Our Corollary 1 also shows that when $K(s,N,\rho)$ is essentially square-integrable, then so is $\gamma\cdot K(s,N,\rho)$, with the same formal degree.
Combining Corollary 1 with the results of [BushKutzBook], [BushKutzComp], [BHK], we obtain the second result of this note: We give a local proof, valid for function fields, for $G=\mathrm{GL}_{n}$, of
Theorem 2 (Clozel, unpublished).
Let $G$ be a connected reductive group over a $p$-adic field $F$, and let $\pi$ be an essentially square-integrable representation of $G$. Then $\gamma\cdot\pi$ is again essentially square-integrable, with the same formal degree as $\pi$.
An exposition of Clozel’s proof is given in [KSV], where the corresponding result for function fields is then deduced from Theorem 2 by Kazhdan’s method of close fields.
Finally, note that irreducible tempered and discrete series representations are standard, but that within the class of standard representations, the tempered representations are not stable under $\mathrm{Aut}(\mathbb{C})$. Thus the standard representations are a class of representations defined over $\bar{\mathbb{Q}}_{\ell}$ containing the tempered representations. Our proof of Theorem 1 will show that the standard representations form essentially the minimal enlargement of the tempered representations, at least in the case of the principal block.
In [BravermanKazhdan], Braverman-Kazhdan introduce a certain subalgebra $\mathcal{J}$ of the Harish-Chandra Schwartz algebra of $G(F)$. By [suzuki], Theorem 1 implies that the class of simple $\mathcal{J}$-modules is defined over $\bar{\mathbb{Q}}_{\ell}$. Our proof in this note provides an independent argument for the principal block.
In future work, we hope to expand our techniques to unipotent blocks of classical groups, using [Kato] and [CiubotaruKato].
1.2 Acknowledgements
The author thanks Anne-Marie Aubert, Dinakar Muthiah, Evgeny Shinder, Maarten Solleveld, Eric Sommers, and Yakov Varshavsky for helpful conversations. Some of this work was completed during “Geometric representation theory in the Langlands program” at CIRM Luminy; the author thanks the organizers and staff for their hospitality. The author was supported by Engineering and Physical Sciences Research Council grant UKRI167 “Geometry of double loop groups.”
2 Standard modules
2.1 Notation and conventions
Let $F$ be a local non-archimedean field and let $G^{\vee}$ be a connected reductive group defined and split over $F$ (we adopt this convention on which group is dual, as the $p$-adic group will appear little in the sequel.) Let $q$ be the cardinality of the residue field of $F$.
Let $G$ be the group dual to $G^{\vee}$, considered as a group over $\mathbb{C}$, with root datum $(X^{*},\Phi,X_{*},\Phi^{\vee})$, and chosen set $\Sigma$ of simple roots. We write $\mathcal{B}$ for the flag variety of $G$, $\tilde{\mathcal{N}}=T^{*}\mathcal{B}$, and $\mathrm{St}=\tilde{\mathcal{N}}\times_{\mathcal{N}}\tilde{\mathcal{N}}$ for the Steinberg variety of $G$, where $\mathcal{N}\subset\mathfrak{g}=\operatorname{\mathrm{Lie}}(G)$ is the nilpotent cone. All of these depend only on $G_{\mathrm{ad}}=G/Z(G)$, a semisimple group of adjoint type. Therefore they are each defined over $\mathbb{Q}$, by Chevalley’s theorem (in fact, over $\mathbb{Z}$). Given a nilpotent element $N\in\mathfrak{g}$, we write $\mathcal{B}_{N}$ for the Springer fibre, and given $g\in G$, we write $X^{g}$ for the fixed points for any $G$-scheme $X$. We will view the above $G$-schemes as $G\times{\mathbb{G}_{\mathrm{m}}}$-scheme, such that $\mathbb{G}_{\mathrm{m}}$} scales the cotangent fibres in $\tilde{\mathcal{N}}$.
We write $K(X)=K_{0}(\mathrm{Coh}(X))$ for the Grothendieck group, for any stack $X$.
[…]
$H\simeq K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})$
as in [CG, §7.6]. When $\mathbf{q}=q$ is the size of the residue field, it is well-known that $H|_{\mathbf{q}=q}=C_{c}(I\backslash G^{\vee}(F)/I,\mathbb{Z})$ is the algebra of $I$-biinvariant compactly supported functions on $G^{\vee}(F)$.
2.2 Action of field automorphisms
Now we fix once and for all a field automorphism $\gamma\in\mathrm{Aut}(\mathbb{C})$. Following [KSV, §2], [SolleveldStandard, §5], let $\mathbb{C}_{\gamma}$ denote the $\mathbb{C},\mathbb{C}$-bimodule $\mathbb{C}$ with actions
$z_{1}\cdot z\cdot z_{2}=z_{1}z\gamma(z_{2}).$
Let $\pi$ be an admissible $G^{\vee}(F)$-module, and let $\gamma\cdot\pi=\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\pi$ be the twist of $\pi$ by $\gamma$. It has scalar multiplication given by
$z\cdot(1\otimes v)=z\otimes v=1\otimes\gamma(z)^{-1}v$
and $G^{\vee}(F)$-action given by
$(\gamma\cdot\pi)g\cdot(1\otimes v)=1\otimes\pi(g)v.$
The matrix coefficients of $\gamma\cdot\pi$ are obtained from the matrix coefficients of $\pi$ by applying $\gamma$. Indeed, clearly we have $\widetilde{\gamma\cdot\pi}=\gamma\cdot\widetilde{\pi}$, with pairing
$\left\langle z_{1}\otimes\tilde{v},z\otimes v\right\rangle_{\gamma\cdot\pi}=z_{1}z\gamma(\left\langle\tilde{v},v\right\rangle_{\pi}).$
This means that for $1\otimes v=1\otimes v_{\alpha_{0}}\in\gamma\cdot\pi$ for $\left\{v_{\alpha}\right\}_{\alpha}$ some basis of $\pi$, and $1\otimes\tilde{v}\in\gamma\cdot\tilde{\pi}$, such that $\left\langle\tilde{v},v\right\rangle_{\pi}=1$, then writing
$(\gamma\cdot\pi)(g)(1\otimes v)=1\otimes\pi(g)v=1\otimes\left(\sum_{\alpha}c_{\alpha}v_{\alpha}\right)=\sum_{\alpha}\gamma(c_{\alpha})\otimes v_{\alpha},$
we have
$\left\langle 1\otimes\tilde{v},(\gamma\cdot\pi)(g)(1\otimes v)\right\rangle_{\gamma\cdot\pi}=\left\langle 1\otimes\tilde{v},1\otimes\pi(g)v\right\rangle=\gamma(c_{\alpha})=\gamma\left(\left\langle\tilde{v},\pi(g)v\right\rangle_{\pi}\right).$
Given $f\in\mathcal{H}=C_{c}^{\infty}(G^{\vee}(F))$, then matrix coefficients of the operator $(\gamma\cdot\pi)(f)$ are then by definition the numbers
$\left\langle 1\otimes\tilde{v},(\gamma\cdot\pi)(f)(1\otimes v)\right\rangle_{\gamma\cdot\pi}=\int{f(g)\left\langle 1\otimes\tilde{v},(\gamma\cdot\pi)(g)(1\otimes v)\right\rangle_{\gamma\cdot\pi}}\,\mathrm{d}{g}=\int{f(g)\gamma\left(\left\langle\tilde{v},\pi(g)v\right\rangle\right)}\,\mathrm{d}{g}.$ (1)
As we normalize the Haar measure such that open compact subgroups of $G$ have rational volumes, if $f\in C_{c}^{\infty}(G,\mathbb{Q})$ is rational-valued, then the above is equal to
$\gamma\left(\left\langle\tilde{v},\pi(f)v\right\rangle\right),~~~f\in C_{c}^{\infty}(G^{\vee}(F),\mathbb{Q}).$
2.3 Iwahori-spherical standard modules as vector spaces
Recall that the standard modules are indexed by triples $(s,N,\rho)$ up to $G(\mathbb{C})$-conjugacy, where $s\in G(\mathbb{C})$ is semisimple, $N\in\mathfrak{g}$ is nilpotent, $\mathrm{Ad}(s)N=qN$, and $\rho$ is an irreducible representation of the component group $\pi_{0}(Z_{G}(N,s))$ appearing in $H_{*}(\mathcal{B}_{N}^{(s,q)})$, where we view $\mathcal{B}_{N}$ as a $G\times{\mathbb{G}_{\mathrm{m}}}$-variety as above. Then the corresponding standard module is by definition, $K(s,N,\rho)=H_{*}(\mathcal{B}_{N}^{s})_{\rho}$, the $\rho$-isotypic part of $H_{*}(\mathcal{B}_{N}^{s})$. This is shown in [KLDeligneLanglands] [CG] under the assumption that the derived subgroup of $G$ is simply-connected, and extended to the general case in [ReederIsogeny, Thm. 3.5.4].
The adjoint representation depends by definition only on the adjoint form $G_{\mathrm{ad}}$ of $G$, which, by Chevalley’s theorem, is certainly defined over $\mathbb{Q}$, as is $\mathfrak{g}$ (see, e.g., [LusJAMS2009]). Therefore the adjoint representation of $G_{\mathrm{ad}}$ on $\mathfrak{g}$ is defined over $\mathbb{Q}$ [HumphreysLAG, §34.2 (9.1), (10.3)], and hence each nilpotent orbit of $G$ is defined over $\mathbb{Q}$ [HumphreysLAG, §34.2 (8.3)]. By the proof of Theorem 4.2 of [Kottwitz], each nilpotent orbit in $\mathfrak{g}$ therefore has a $\mathbb{Q}$-point, so we may assume that $N$ is $\gamma$-fixed. Therefore we have
$(\mathcal{B}_{N}^{s})_{\gamma}=\mathcal{B}_{N}^{s}\times_{\operatorname{Spec}\mathbb{C}}\operatorname{Spec}\mathbb{C}=\mathcal{B}_{N}^{\gamma(s)}$$\operatorname{Spec}\mathbb{C}$$\mathbb{C}$$\mathcal{B}_{N}^{s}$$\operatorname{Spec}\mathbb{C}$$\mathbb{C}$$\scriptstyle{\operatorname{Spec}(\gamma)}$$\scriptstyle{\operatorname{Spec}(\gamma)}$$\scriptstyle{\gamma}$ (2)
where we abusively write $\operatorname{Spec}(\gamma)$ for both the original morphism and its base-change, a convention we will use throughout when there is no danger of confusion.
Diagram (2) exhibits $\mathcal{B}_{N}^{s}$ and $\mathcal{B}_{N}^{\gamma(s)}$ as isomorphic $\mathbb{Q}$-varieties, but the two are not isomorphic $\mathbb{C}$-varieties unless $\gamma=\mathrm{id}_{\mathbb{C}}$, and $\gamma$ does not even induce a continuous map of analytifications unless it is complex conjugation or the identity. None the less, there is still a natural map on homology, thanks to
Theorem 3 ([DLP]).
For any $(s,N)$ such that $\mathrm{Ad}(s)N=qN$, we have an isomorphism of Chow group and homology
$A_{*}(\mathcal{B}_{N}^{s})\overset{\sim}{\to}H_{*}(\mathcal{B}_{N}^{s};\mathbb{Z})$
via the cycle class map.
Writing $A_{*}(X)_{\mathbb{Q}}=A_{*}(X)\otimes_{\mathbb{Z}}\mathbb{Q}$ for any scheme $X$, we therefore consider the standard modules as realized on $\mathbb{C}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$, and and define a $\mathbb{Q}$-linear isomorphism
$\operatorname{Spec}(\gamma)^{*}\colon A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}\overset{\sim}{\to}A_{*}(\mathcal{B}_{N}^{\gamma(s)})_{\mathbb{Q}}$ (3)
by flat pullback of algebraic cycles. We will occasionally refer to [CG] for results written in the language of Borel-Moore homology, this is harmless by Theorem 3. Note that by [Fulton, Cor. 18.3.2], the Chern character
$\mathrm{ch}_{*}\colon K(\mathcal{B}_{N}^{s})\otimes_{\mathbb{Z}}\mathbb{Q}\overset{\sim}{\to}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$
is an isomorphism.
2.4 Standard modules as modules over the Bernstein subalgebra
We define an action of $G$-equivariant line bundles on the diagonal $\tilde{\mathcal{N}}\subset\mathrm{St}$ on $A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$: Let $\mathcal{L}$ be such a line bundle, and consider its class as belonging to $K(\mathcal{B})$ via the Thom isomorphism. Consider the natural embeddings
$\mathcal{B}^{s}$$\mathcal{B}$$\mathcal{B}^{\gamma(s)}=(\mathcal{B}^{s})_{\gamma}$$\mathcal{B}$$\scriptstyle{\iota}$$\scriptstyle{\operatorname{Spec}(\gamma)}$$\scriptstyle{\iota_{\gamma}}$$\scriptstyle{\operatorname{Spec}(\gamma)}$
and let $v\in A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$. Then
$\mathcal{L}\cdot v:=\mathrm{ch}^{*}(\iota^{*}\mathcal{F})\frown v$
is the action of cohomology on Borel-Moore homology [CG, §2.6.40] and $\mathrm{ch}^{*}$ is the cohomological Chern character [BFMIHES] [CG, §5.8].
Lemma 1.
The $\mathbb{Q}$-linear isomorphism $\operatorname{Spec}(\gamma)^{*}$ in (3) is linear for the $\cdot$-action of line bundles.
Proof.
This is entirely formal but notationally involved. Write $\operatorname{Spec}(\gamma)$ for all the vertical maps in the pair of pullback squares
$(\mathcal{B}_{N}^{s})_{\gamma}$$(\mathcal{B}^{s})_{\gamma}$$\operatorname{Spec}\mathbb{C}$$\mathcal{B}_{N}^{s}$$\mathcal{B}^{s}$$\operatorname{Spec}\mathbb{C}.$}
Let $\mathcal{E}\in K(\mathcal{B}_{N}^{s})$ be such that $\mathrm{ch}_{*}(\mathcal{E})=v$.
We have, for any line bundle $\mathcal{L}$ on $\Delta\tilde{\mathcal{N}}$,
$\operatorname{Spec}(\gamma)^{*}(\mathcal{L}\cdot v)$ $=\operatorname{Spec}(\gamma)^{*}(\mathrm{ch}^{*}(\iota^{*}\mathcal{L})\frown\mathrm{ch}_{*}(\mathcal{E}))$ (4)
$=\operatorname{Spec}(\gamma)^{*}(\mathrm{ch}_{*}(\iota^{*}\mathcal{L}\otimes\mathcal{E})$ (5)
$=\mathrm{ch}_{*}(\operatorname{Spec}(\gamma)^{*}\iota^{*}\mathcal{L}\otimes\operatorname{Spec}(\gamma)^{*}\mathcal{E})$ (6)
$=\mathrm{ch}_{*}(\iota_{\gamma}^{*}\operatorname{Spec}(\gamma)^{*}\mathcal{L}\otimes\operatorname{Spec}(\gamma)^{*}\mathcal{E})$
$=\mathrm{ch}^{*}(\iota_{\gamma}^{*}\operatorname{Spec}(\gamma)^{*}\mathcal{L})\frown\mathrm{ch}_{*}(\operatorname{Spec}(\gamma)^{*}\mathcal{E})$ (7)
$=\mathrm{ch}^{*}(\iota_{\gamma}^{*}\mathcal{L})\frown\operatorname{Spec}(\gamma)^{*}v,$ (8)
where between (4) and (5) we used smoothness of $\mathcal{B}^{s}$ and [Fulton, Prop. 18.1 (c)], and likewise between (7) and (8); between (5) and (6) and (7) and (8), we also used that the homological Chern character commutes with pullback by smooth proper morphisms [BFMIHES, (2.6)], [Fulton, Thm. 18.1]. Between (7) and 8 we used that $\operatorname{Spec}(\gamma)^{*}[\mathcal{L}]=[\mathcal{L}]$ in the Grothendieck group, as the associated bundle construction defining each generator $\mathcal{O}_{\lambda}$ of $K(\tilde{\mathcal{N}}/G\times{\mathbb{G}_{\mathrm{m}}})$, $\lambda\in X^{*}(T)$, is defined over $\mathbb{Q}$ (in particular this implies the equality in equivariant $K$-theory, see also Step 1 of the proof of Proposition 1). Between (5) and (6) we used that $\operatorname{Spec}(\gamma)$ is flat and $\iota^{*}\mathcal{L}$ is a line bundle. ∎
Now we recall the definition of the $H$-module structure on $A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{C}}$ from [CG, §8.2]. Chose a reference Borel subgroup $B_{0}$ of $G_{\mathrm{ad}}$ such that $\mathcal{B}=G_{\mathrm{ad}}/B_{0}$. We may chose $B_{0}$ to be defined over $\mathbb{Q}$, so that it is $\gamma$-stable. Let $B$ be another Borel subgroup of $G_{\mathrm{ad}}$, so that there is $g\in G(\mathbb{C})$ such that $B=gB_{0}g^{-1}$. By [CG, 6.1.1] the induced isomorphism $\varphi_{B}$
[…]
of tori, depends only on $B$, not on $g$. Now for $\mathcal{O}_{\lambda}$ on $\Delta\tilde{\mathcal{N}}\subset\mathrm{St}$, the action of $\mathcal{O}(\lambda)$ on $\mathbb{C}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$ is defined in [CG, §8.2] as follows: Let $v$ be a linear combination of cycles contained in a single connected component of $\mathcal{B}_{N}^{s}$. This component is contained in a unique connected component $\mathcal{B}_{j}$ of $\mathcal{B}^{s}$, and $\mathcal{B}_{j}=Z_{G}(s)\cdot B_{j}$ for some Borel subgroup $B_{j}$ [CG, 8.8.7]. As $s\in B_{j}$ and other Borel subalgebras in $\mathcal{B}_{j}$ are $Z_{G}(s)$-conjugate, it is well-defined to put
$\mathcal{O}(\lambda)\star(1\otimes v):=\lambda(\varphi_{B_{j}}(s))\otimes(\mathcal{O}(\lambda)\cdot v).$
The action is then extended to the entire Chow group by linearity.
Definition 1.
Let $\Phi(\gamma)$ be the isomorphism of left $\mathbb{C}$-modules
$\Phi(\gamma)\colon\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}\right)\to\mathbb{C}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{\gamma(s)})_{\mathbb{Q}}$
given by
$\left(z_{1}\otimes\left(z_{2}\otimes v\right)\right)\mapsto z_{1}\gamma(z_{2})\otimes\operatorname{Spec}(\gamma)^{*}v.$
The automorphism $\gamma$ induces an isomorphism
$\gamma\colon Z_{G}(s,N)\to Z_{G}(\gamma(s),N),$ (10)
and hence an isomorphism of component groups $\pi_{0}(Z(s,N))\simeq\pi_{0}(Z(\gamma(s),N))$, such that the diagram
$\mathcal{B}_{N}^{\gamma(s)}$$\mathcal{B}_{N}^{\gamma(s)}$$\operatorname{Spec}(\mathbb{C})$$\mathcal{B}_{N}^{s}$$\mathcal{B}_{N}^{s}$$\operatorname{Spec}\mathbb{C}$$\scriptstyle{\gamma(g)}$$\scriptstyle{\operatorname{Spec}(\gamma)}$$\scriptstyle{\operatorname{Spec}(\gamma)}$$\scriptstyle{\operatorname{Spec}(\gamma)}$$\scriptstyle{g}$ (11)
commutes.
Lemma 2.
The map $\Phi(\gamma)$ is $\pi_{0}(Z_{G}(N,s))$-linear, upon identifying component groups via (10).
Proof.
This is exactly commutativity of (11). ∎
Lemma 3.
The map $\Phi(\gamma)$ intertwines, for each $\mathcal{O}\lambda$ the actions
$\mathcal{O}\lambda(z_{1}\otimes(z_{2}\otimes v))=z_{1}\otimes(\mathcal{O}_{\lambda}\star(z_{2}\otimes v))$
[…]
Proof.
If $B_{j}=gB_{0}g^{-1}$ is in the connected component of $\mathcal{B}^{s}$ containing the component of $\mathcal{B}_{N}^{s}$ in which a cycle $v$ is contained, then $B_{\gamma(j)}:=\gamma(g)\gamma(B_{0})\gamma(g)^{-1}=\gamma(g)B_{0}\gamma(g)^{-1}$ is in the connected component of $\mathcal{B}^{\gamma(s)}$ in which $\gamma^{*}v$ is contained. Hence we have
$\lambda(\varphi_{B_{\gamma(j)}}(\gamma(s)))=\lambda(\gamma(g)\gamma(s)\gamma(g)^{-1})=\gamma(\lambda(gsg^{-1}))=\gamma(\lambda(\varphi_{B}(s))$
as required. The Lemma is true for general homology classes by linearity. ∎
2.5 Standard and essentially square-integrable representations
Our first main result is the following theorem and its corollary.
Theorem 4.
The map $\Phi(\gamma)$ is $H$-linear, and commutes with the action of the component group of $Z_{G}(s,N)\simeq Z_{G}(\gamma(s),N)$. Hence we obtain an isomorphism
$\Phi(\gamma)\colon\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}K(s,N,\rho)\overset{\sim}{\to}K(\gamma(s),N,\rho),$
$H$-modules, where the action on the left hand side is via
$h\cdot(z_{1}\otimes v)=z_{1}\otimes h\cdot v.$
Proof.
Combine Proposition 1 below and Lemma 2. ∎
Remark 1.
For generic $s$ and $\gamma(s)$, the standard modules $K(s,N,\rho)$ and $K(\gamma(s),N,\rho)$ are simple. By [EvensMirkovic, Thm. 5.5], Lemma 3 already implies that $K(s,N,\rho)\simeq K(\gamma(s),N,\rho)$ in this case.
Definition 2.
A smooth admissible representation of $\omega$ of $G^{\vee}(F)$ is square-integrable modulo the centre $Z(G^{\vee}(F))$ if $\omega$ admits a unitary central character and all matrix coefficients of $\omega$ are square-integrable functions on $G^{\vee}(F)/Z(G(F))$.
A smooth admissible representation $\omega$ of $G^{\vee}(F)$ is essentially square integrable if it is of the form $\omega\otimes\nu$ for an unramified character $\nu$ and square-integrable modulo centre representation $\omega$ of $G^{\vee}(F)$.
Recall that standard representation $K(s,N,\rho)$ is an irreducible essentially square-integrable representation, if and only if, there is no proper parabolic subgroup of $G$ with Levi subgroup containing $(s,u)$, where $u=\exp(N)$ [KLDeligneLanglands, Thm. 8.3 (a)].
Corollary 1.
Let $\pi$ be an essentially square-integrable representation of $H$. Then $\gamma\cdot\pi$ is essentially square-integrable, and we have an equality of formal degrees
$d(K(s,N,\rho)=d(K(\gamma(s),N,\rho).$
More generally, if $\pi$ is a standard $H$-module, so is $\gamma\cdot\pi$, and twist by $\gamma$ also preserves genericity of the Langlands quotients.
Proof.
The last statement is a restatement of Theorem 4.
If $\pi$ is square-integrable, then $\pi=K(s,N,\rho)$ for some $\rho$. Clearly, for $u$ as above, $(s,u)$ is not contained in any proper Levi subgroup of $G$ if and only if same is true for $(\gamma(s),u)$. Therefore the last statement follows from Theorem 4.
Now we argue for the equality of formal degrees. Recall from [Reeder, Formula A], that we have
$d(K(\gamma(s),N,\mathrm{triv}))^{-1}=q^{\dim\mathcal{B}}\sum_{J\subset\Sigma}\frac{1}{P_{J}(q)}\sum_{\lambda\in\Lambda_{J}}|M(\lambda,\gamma(s))|^{2},$
where $W_{J}$ is the parabolic subgroup of $W$ generated by $J\subset\Sigma$, $P_{J}(q)$ is its Poincaré polynomial,
$\Lambda_{J}=\left\{\lambda\in X^{*,+}\,\middle|\,\mathrm{Stab}_{W}(\lambda)=W_{J}\right\},$
where $X^{*,+}$ is the set of dominant weights, and $M(\lambda,s)\colon\Lambda_{J}\to\mathbb{C}$ is the function
$M(\lambda,s)=\sum_{w\in W}w\left(\lambda\prod_{\alpha>0}\frac{1-q^{-1}\alpha}{1-\alpha}\right)(s).$
The product is taken over all positive roots, and the action of the Weyl group is the natural one on rational functions on $T$.
Recall that from $(s,N)$ $q$-commuting we can obtain commuting $(s_{1},N)$: Let $\phi$ be a homomorphism
$\phi\colon\mathrm{SL}_{2}\to G$
obtained from $N$ via the Jacobson-Morozov theorem. Then $s=s_{1}\phi(\mathrm{diag}(q^{-1/2},q^{1/2}))$ and $\mathrm{Ad}(s_{1})N=N$.
Now, in general, $\gamma$ of course does not commute with complex conjugation. In the case at hand, though, square-integrability of $K(s,N,\rho)$ and $K(\gamma(s),N,\rho)$ implies that $s_{1}$ and $\gamma(s_{1})$ belongs to the maximal compact subgroup of $T^{\vee}$. Therefore
$\lambda(\gamma(s))=\lambda(\gamma(s_{1}))\lambda(\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2}))$
with $\lambda(\gamma(s_{1}))\in U(1)\subset\mathbb{C}^{\times}$ for any weight $\lambda$, and, by hypothesis, likewise for $s_{1}$ (note $\gamma(s_{1})=\gamma(s)\phi(\gamma(q^{-1/2},q^{1/2})=\gamma(s)\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2})$). Therefore
$\overline{\lambda(w\gamma(s))}=\lambda(w\gamma(s_{1}))^{-1}\lambda(\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2}))\\ =\lambda(w\gamma(s_{1}^{-1}))\lambda(\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2}))=\gamma\left(\lambda(ws_{1}^{-1})\right)\lambda(\phi(\pm\mathrm{diag}(q^{-1/2},q^{1/2}))=\gamma\left(\overline{\lambda(ws)}\right).$
It follows that
$|M(\lambda,\gamma(s))|^{2}=M(\lambda,\gamma(s))\overline{M(\lambda,\gamma(s))}=\gamma\left(|M(\lambda,s)|^{2}\right),$
and so $d(K(\gamma(s),N,\mathrm{triv}))=\gamma(d(K(s,N,\mathrm{triv}))$. On on the other hand, by [FOS], $d(K(s,N,\rho))\in\mathbb{Q}$. For general $\rho\neq\mathrm{triv}$, we have
$d(K(s,N,\rho))=\dim(\rho)d(K(s,N,\mathrm{triv}))$
by op. cit.
3 Essentially discrete series representations of $\mathrm{GL}_{n}$
3.1 Types and Hecke algebras
In this section, we give our proof of Theorem 2 for $G=\mathrm{GL}_{n}$. We first recall
Theorem 5 ([BushKutzBook], [BushKutzComp]).
Let $F$ be a non-archimedean local field and let $G=\mathrm{GL}_{n}(F)$; let $q$ be the cardinality of the residue field of $F$. For any Bernstein block $\mathbf{Rep}_{\mathfrak{s}}(G)$ of the category of smooth representations $\mathbf{Rep}(G)$ of $G$, there are integers $\left\{m_{i}\right\}_{i}$ and $\left\{f_{i}\right\}_{i}$ and an equivalence of categories
$\mathbf{Rep}_{\mathfrak{s}}(G)\simeq\bigotimes_{i}H_{m_{i}}(q^{f_{i}}),$
for integers $m_{i}$ and $f_{i}$, where $H_{m_{i}}(q^{f_{i}})$ is an affine Hecke algebra of type $\tilde{A}_{m_{i}}$ with parameter $q^{f_{i}}$, and the tensor product is taken over $\mathbb{C}$.
Theorem 6 ([BHK], Thm. 4.3).
The equivalences of Theorem 5 preserve Plancherel measures up to a constant factor, and hence preserve discrete series and tempered representations.
Note that Theorem 6 applies to any type, not only the ones constructed in [BushKutzBook], [BushKutzComp].
As remarked in the introduction, if $\sigma$ is a supercuspidal representation of $G$, then clearly $\gamma\cdot\sigma$ is again supercuspidal. The compatibility with normalized parabolic induction, up to unramified twist of the induced representation, of $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}(-)$ [KSV, (5.12)] then implies that the auto-equivalence $\pi\mapsto\gamma\cdot\pi$ of $\mathbf{Rep}(G)$ sends the Bernstein block for an inertial class $\mathfrak{s}=[M,\sigma]_{G}$ to the Bernstein block for $\mathfrak{s}_{\gamma}=[M,\gamma\cdot\sigma]_{G}$. As remarked above, $T,\mathrm{triv}]_{G}=[T,\gamma\cdot\mathrm{triv}]_{G}$, but in general $\mathfrak{s}_{\gamma}$ labels a new block. If $(K,\rho)$ is a type for $\mathfrak{s}$, then $(K,\gamma\cdot\rho)$ is a type for $\mathfrak{s}_{\gamma}$, and we have an isomorphism of rings
$H(G,K,\rho)\to H(G,K,\gamma\cdot\rho)$
given by
$f\mapsto(g\mapsto\gamma(f(g)),$
which is $\gamma$-semilinear, where $\gamma(f(g))$ means we apply $\gamma$ to the matrix of coefficients of $f(g)\in\operatorname{End}_{\mathbb{C}}(\rho)$. Equivalently, we have a $\mathbb{C}$-linear isomorphism
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}H(G,K,\rho)\to H(G,K,\gamma\cdot\rho)$
given by
$z\otimes f\mapsto z\gamma(f).$
Rewriting this in terms of affine Hecke algebras, we have isomorphisms
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\bigotimes_{i}H_{m_{i}}(q^{f_{i}})\overset{\sim}{\to}H(G,K,\gamma\cdot\rho)\simeq\bigotimes_{i}H_{m_{i}}(q^{f_{i}}),$
where the second isomorphism is induced by the auto-equivalence $\pi\mapsto\gamma\cdot\pi$.
Now suppose that $\pi$ is a simple essentially square-integrable representation in $\mathbf{Rep}(G)_{\mathfrak{s}}$. Then $\pi$ corresponds to a $H(G,K,\rho)$-module
$\bigotimes_{i}K(s_{i},N_{i})$ (12)
and $\gamma\cdot\pi$ corresponds to the $H(G,K,\gamma\cdot\rho)$-module
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\bigotimes_{i}K(s_{i},N_{i}),$ (13)
by the discussion in Section 2. Now we give our proof of Theorem 2.
3.2 Proof of Theorem 2 for $G=\mathrm{GL}_{n}$
To show that $\gamma\cdot\pi$ is again essentially discrete series, it suffices by Section 3.1 to show that the $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\bigotimes_{i}H_{m_{i}}(q^{f_{i}})$-module (13) is isomorphic via $\bigotimes_{i}\Phi(\gamma)$ to $\bigotimes_{i}K(\gamma(s_{i}),N_{i})$ under the second isomorphism above to $\bigotimes_{i}H_{m_{i}}(q^{f_{i}})$. This also implies the statement about formal degrees, as formal degrees of essentially discrete series representations of affine Hecke algebras of type $\mathrm{GL}_{n}$ depend only on $N$. Therefore, it suffices to prove
Proposition 1.
Let $H$ be an affine Hecke algebra. Then the diagram
$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})_{\mathbb{Q}}$$\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{pt}/\left\langle s\right\rangle)\otimes_{\mathbb{Q}}K(\mathrm{St}^{(s,q)})_{\mathbb{Q}}$$\mathbb{C}\otimes_{\mathbb{Q}}K(\mathrm{pt}/\left\langle\gamma(s)\right\rangle)\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q)})$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}_{a}\otimes_{\mathbb{Q}}K(\mathrm{St}^{s,q})_{\mathbb{Q}}\right)$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\left(\mathbb{C}_{a}\otimes_{\mathbb{Q}}K(\mathrm{St}^{s,q})_{\mathbb{Q}}\right)$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}K(\mathrm{St}^{\gamma(s),q})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}_{s}\otimes_{\mathbb{Q}}A_{*}(\mathrm{St}^{s})_{\mathbb{Q}}$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}A_{*}(\mathrm{St}^{\gamma(s)})_{\mathbb{Q}}$$\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\mathbb{C}_{s}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$$\mathbb{C}_{\gamma(s)}\otimes_{\mathbb{Q}}A_{*}(\mathcal{B}_{N}^{\gamma(s)})_{\mathbb{Q}},$$\scriptstyle{i_{s}^{*}\circ\mathrm{Forg}_{\left\langle s\right\rangle}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{i_{\gamma(s)}\circ\mathrm{Forg}_{\left\langle\gamma(s)\right\rangle}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{ev}}$$\scriptstyle{\mathrm{ev}}$$\scriptstyle{\mathrm{id}\otimes(-\otimes(1\boxtimes\lambda_{s}^{-1}))}$$\scriptstyle{(\mathrm{id}\cdot\gamma)\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes(-\otimes(1\boxtimes\lambda_{\gamma}(s)^{-1}))}$$\scriptstyle{(\mathrm{id}\cdot\gamma)\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{RR}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{RR}}$$\scriptstyle{\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{act}(v)}$$\scriptstyle{\mathrm{id}\otimes\mathrm{id}\otimes\mathrm{act}(\Phi(\gamma)v)}$$\scriptstyle{\Phi(\gamma)}$
where the columns encode the action of a single $H_{m_{i}}(q^{f_{i}})$ on a single $K(s_{i},N_{i})$ as per [CG, §8.2], commutes. (Here, we have written $q=q^{f_{i}}$, $s=s_{i}$, etc. to unburden notation.)
In the diagram, we write $\left\langle s\right\rangle$ for the smallest closed subgroup of $G$ containing $s$. Note also that that $H_{*}(\mathrm{St}^{s};\mathbb{Q})$ and $H_{*}(\mathrm{St}^{\gamma(s)};\mathbb{Q})$ are isomorphic to $A_{*}(\mathrm{St}^{s})_{\mathbb{Q}}$ and $A_{*}(\mathrm{St}^{\gamma(s)})_{\mathbb{Q}}$, respectively. This follows from [Fulton, Cor. 18.3.2] and [CG, Thm. 6.2.4] (the proof in loc. cit. works for any field coefficients). The Künneth formula for homology and [CG, (5.2.4)] reduce the general case to products of the case of the Proposition.
Proof.
We explain the meaning and commutativity of each box in the diagram. Modulo notation, the proof is almost entirely formal: at least on the tensor factors involving $K$-theory or Chow groups, it essentially amounts to various constructions commuting with isomorphisms of schemes, i.e., being well-defined.
It is enough to check commutativity of the square on elements of the form $1\otimes 1\otimes\mathcal{F}$ for generators $\mathcal{F}$ of $K(\mathrm{St}/G\times{\mathbb{G}_{\mathrm{m}}})$. The bundles $\mathcal{O}_{\lambda}$ on the diagonal component $\Delta\tilde{\mathcal{N}}\subset\mathrm{St}$, pulled back from the line bundles $\mathcal{O}(\lambda)$ on $\mathcal{B}$ generate the Bernstein subalgebra. We checked commutativity for the Bernstein subalgebra in Section 2.4.
We may take the following classes as generators of the finite Hecke algebra $H_{W}$, following [CG, (7.6.1)]: let $\bar{Y}_{\sigma}\subset\mathcal{B}\times\mathcal{B}$ be the Schubert variety indexed by the simple reflection $\sigma$ in $W$. The projection $\bar{Y}_{\sigma}\to\mathcal{B}$ to the first coordinate is smooth with $\mathbb{P}^{1}$ fibres, hence the relative cotangent bundle is a line bundle on $\bar{Y}_{\sigma}$, namely $\Omega^{1}_{\bar{Y}_{\sigma}/\mathcal{B}}=\operatorname{Ind}(P_{\alpha}/B\times^{B}\mathbb{C}_{-\alpha})$, where $\operatorname{Ind}\colon\mathrm{Coh}(P_{\alpha}/B)\to\mathrm{Coh}(\bar{Y}_{\sigma}/G)$ is the induction functor of [CG, §5.2.16] and $\alpha$ is the simple root corresponding to $\sigma$. Finally, our generator is $\mathcal{L}_{\sigma}:=\pi_{\sigma}^{*}\Omega^{1}_{\bar{Y}_{\sigma}/\mathcal{B}}$, where $\pi_{\sigma}\colon T_{\bar{Y}_{\sigma}}^{*}(\mathcal{B}\times\mathcal{B})\to\bar{Y}_{\sigma}$ is the conormal bundle projection.
Example 1.
If $G=\mathrm{SL}_{2}$, then $\mathcal{L}_{\sigma}=\mathcal{O}\boxtimes\mathcal{O}(-2)$ on the zero section component $\mathbb{P}^{1}\times\mathbb{P}^{1}$ of $\mathrm{St}$.
Now we check commutativity.
  1. 1.
    We check the topmost square, which subsumes the first three isomorphisms in [CG, (8.1.6)]. The morphisms $i_{s}$ and $i_{\gamma(s)}$ are inclusion of $s$-, respectively $\gamma(s)$-fixed points, and the forgetful maps are induced by the corresponding obvious functors.
    Taking a generator $\mathcal{L}_{\sigma}$, the character that appears in the $K(\mathrm{pt}/\left\langle s\right\rangle)$-coordinate upon applying $i_{s}^{*}\circ\mathrm{Forg}_{\left\langle s\right\rangle}$ is of the form $\alpha(\varphi_{B_{j}}(s))$ as in (9), where $\mathcal{B}_{j}$ is a connected component of $\mathcal{B}^{s}$ the proof of Lemma 3. We conclude again that the square commutes. Finally, note that $\operatorname{Spec}(\gamma)^{*}[\mathcal{L}_{\sigma}]=\mathcal{L}_{\sigma}$, as the Schubert varieties and their conormal bundles are defined over $\mathbb{Q}$. This holds as equivariant sheaves: concretely, the equivariant structure is a vector bundle morphism $G\times\operatorname{\underline{Spec}}{\operatorname{Sym}\mathcal{L}_{\sigma}^{\vee}}\to\operatorname{\underline{Spec}}{\operatorname{Sym}\mathcal{L}_{\sigma}^{\vee}}$ over $T^{*}_{\bar{Y}_{\sigma}}(\mathcal{B}\times\mathcal{B})$ commuting with $G$ action and fibrewise linear, and this morphism is defined over $\mathbb{Q}$. Therefore this square commutes.
  2. 2.
    Now we check the second square, where, e.g., the top horizontal map sends $\alpha\otimes(z\otimes f\otimes\mathcal{F})\mapsto\alpha\gamma(z)(f\circ\gamma)\otimes\operatorname{Spec}(\gamma)^{*}\mathcal{F}$. It clearly commutes and all the maps in the diagram are isomorphisms.
  3. 3.
    We explain why the third square commutes. Recall from [CG, §5.11] that
    $\lambda_{s}:=\bigotimes_{\alpha\in\mathrm{Sp}(\tilde{\mathcal{N}})}\sum_{i}(-\alpha(s))^{i}\Lambda^{i}N_{\alpha},$
    where $N=\bigoplus_{\alpha\in\mathrm{sp}(N)}N_{\alpha}$ is the $\left\langle s\right\rangle$-weight decomposition of the normal bundle $N$ to $\tilde{\mathcal{N}}^{s}$, and that, as remarked in loc. cit.,
    $\mathcal{O}_{\tilde{\mathcal{N}}}\boxtimes\lambda_{s}^{-1}$
    exists in complexified $K$-theory [CG, Cor. 5.11.3]. It is then clear that $\mathrm{id}\cdot\gamma\otimes\operatorname{Spec}(\gamma)^{*}\lambda_{s}=\lambda_{\gamma(s)}$, and hence the same is true for the inverse classes. Here, we again use that the considerations of (9) apply also the bundles $\mathcal{L}_{\sigma}$, by their construction recalled above. Therefore this square commutes.
  4. 4.
    For the fourth square, note that the bivariant Riemann-Roch map
    $\mathrm{RR}\colon K_{0}(\mathrm{St}^{(s,q)})_{\mathbb{Q}}\overset{\sim}{\to}A_{*}(\mathrm{St}^{(s,q)})_{\mathbb{Q}}$
[…]
    $\mathrm{RR}(\operatorname{Spec}(\gamma)^{*}\mathcal{F})=\operatorname{Spec}(\gamma)^{*}\mathrm{RR}(\mathcal{F})$
    for $\mathcal{F}\in K_{0}(\mathrm{St}^{(s,q)})_{\mathbb{Q}}$; as $\operatorname{Spec}(\gamma)$ is an isomorphism of quasi-projective schemes, its relative Todd class (see B.7.6 of op. cit.) is trivial. Therefore the fourth square commutes.
  5. 5.
    Finally we check the fifth square. Let $\xi$ be an algebraic cycle in $\mathrm{St}^{s}$, and $v$ a cycle in $\mathcal{B}_{N}^{s}$. The right-down circuit gives, for $1\otimes 1\otimes\xi$ and $z_{1}\otimes z_{2}\otimes v$,
    $z_{1}\gamma(z_{2})\otimes\operatorname{Spec}(\gamma)_{\mathcal{B}_{N}^{s}}^{*}p_{1*}(\xi\cap p_{2}^{*}v),$ (14)
    where $p_{i}$ are the respective projections $\mathrm{St}^{s}\to\tilde{\mathcal{N}}^{s}$ and we view $v$ as a cycle in the (smooth, by [CG, Lem. 5.11.1] variety $\tilde{\mathcal{N}}^{s}$ (to which Theorem 6.2.4 of op. cit. also applies). Now we apply base-change [Fulton, Prop. 1.7] (see [CG, §2.7] for the corresponding statements for homology throughout) to the diagram
    $\mathrm{St}^{\gamma(s)}$$\mathrm{St}^{s}$$\tilde{\mathcal{N}}^{\gamma(s)}$$\tilde{\mathcal{N}}^{s}$$\scriptstyle{p_{1}}$$\scriptstyle{\operatorname{Spec}(\gamma)_{\mathrm{St}^{s}}}$$\scriptstyle{p_{1}}$$\scriptstyle{\operatorname{Spec}(\gamma)}$
    noting that $p_{1}$ is proper and the isomorphism $\operatorname{Spec}(\gamma)$ of schemes is certainly flat. Therefore (14) becomes
    $z_{1}\gamma(z_{2})\otimes p_{1*}\left(\operatorname{Spec}(\gamma)^{*}_{\mathrm{St}^{s}}\left(\xi\cap p_{2*}v\right)\right)$ $=z_{1}\gamma(z_{2})\otimes p_{1*}\left(\operatorname{Spec}(\gamma)_{\mathrm{St}^{s}}^{*}\xi\cap\operatorname{Spec}(\gamma)_{\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}}^{*}p_{2}^{*}v\right)$
    $=z_{1}\gamma(z_{2})\otimes p_{1*}\left(\operatorname{Spec}(\gamma)_{\mathrm{St}^{s}}^{*}\xi\cap p_{2}^{*}\operatorname{Spec}_{\tilde{\mathcal{N}}^{s}}(\gamma)^{*}v\right)$
    Here, the first equality used that the intersection pairing commutes with pullback by the vertical flat maps in the diagram
    $\gamma(\xi)$$\tilde{\mathcal{N}}^{\gamma(s)}\times\tilde{\mathcal{N}}^{\gamma(s)}$$\xi$$\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}$$\scriptstyle{\operatorname{Spec}(\gamma)_{\mathrm{St}^{s}}}$$\scriptstyle{\operatorname{Spec}(\gamma)_{\tilde{\mathcal{N}}^{s}\times\tilde{\mathcal{N}}^{s}}}$
    by [Fulton, Thm. 6.2 (b)]. The second equality follows from commutativity of the obvious square, viewing $v$ as a cycle in the ambient smooth variety $\tilde{\mathcal{N}}^{s}$.
The proof of Theorem 2 for $G=\mathrm{GL}_{n}$ is now complete. ∎
Remark 2.
In general, the action of $\mathcal{L}_{\sigma}$ on $H_{*}(\mathcal{B}_{N}^{s};\mathbb{C})$ depends only on the conjugacy class of Cartan subgroup of (the Levi quotient of) the disconnected group $Z_{G}(N)$, see [CellsIV, 3.4 (a)], or [rigid, Lem. 14, §3.5] for this observation in the language of the operator Paley-Wiener theorem for Hecke algebras. In either perspective, either because $Z_{\mathrm{GL}_{n}}(N)$ is connected, or because the only $I$-spherical discrete series of each Levi subgroup of $\mathrm{GL}_{n}$ is the tensor product of Steinberg representations, respectively, it follows that the action depends only on $N$ in this case.